# Decomposing Centrally Symmetric Convex Polyhedral Surfaces into Parallelograms

### 22 Mar 2026 v1

@@ 4 Polyhedra with $8$ Vertices
Therefore, $\psi_{1}$ is continuous and invertible. Since $\psi_{1}$ is a composition of $F_{1}$ by another map, $F_{1}$ is injective on $(0,\infty)^{6}$. By Theorem 2.1, $\mathcal{C}_{8}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})$ has real dimension $6$. By the invariance of domain, $\mathcal{O}_{1}=F_{1}((0,\infty)^{6})$ is open in $\mathcal{C}_{8}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})$. ∎
@@
Given a space $\mathcal{O}$ arising from a loop arrangement $\mathscr{A}$, we can consider the space of polyhedral surfaces in $\mathcal{O}$ with surface area $1$. We call this space the moduli space arising from $\mathscr{A}$, and denote it by $\Delta$. Let $\overline{\Delta}\subseteq\overline{\mathcal{M}_{8}}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})$ be the metric completion of $\Delta$ with respect to the real hyperbolic metric in Theorem 2.2. We denote a face of $\overline{\Delta}$ by $\sigma_{i}({1})$ (where $1$ refers to the surface area) if it is a subset of the face $\sigma_{i}$ in $\overline{\mathcal{O}}$. If $\mathscr{A}^{\prime}$ is another loop arrangement that differs from $\mathscr{A}$ by $l_{i}$ only, we say that $\overline{\Delta}$ and $\overline{\Delta^{\prime}}$ are on different sides of $\bm{\sigma_{i}({1})}$ if and only if $\overline{\mathcal{O}}$ and $\overline{\mathcal{O}^{\prime}}$ are on different sides of $\sigma_{i}$.
@@
The space $\overline{\Delta_{1}}$ has the structure of a real hyperbolic regular ideal $5$-simplex.
@@
Combinatorically, $\overline{\Delta_{1}}$ is a polytope with six vertices and six facets. Each vertex of $\overline{\Delta_{1}}$ has five vanishing coordinates among six with respect to $\mathscr{A}_{1}$. This corresponds to the case where the surface degenerates to a line. Note that each vertex of $\overline{\Delta_{1}}$ belongs to five facets of $\overline{\Delta_{1}}$, and each facet contains five vertices. Therefore, $\overline{\Delta_{1}}$ is a $5$-simplex by Proposition 2.16 in [13].
@@
Similar to $\mathscr{A}_{2}$, by permuting the the labels of vertices and loops in $\mathscr{A}_{1}$, we can obtain more loop arrangements that are different from but combinatorically equivalent to $\mathscr{A}_{1}$. It is clear that there are finitely many such loop arrangements. We consider the metric completions of the moduli spaces arising from these loop arrangements. For convenience, we will call them “$5$-simplices” due to Lemma 4.3. Later, we are going to construct infinitely many copies of such $5$-simplices and regard them as distinct spaces, though some of them arise from the same loop arrangement. Our goal is to glue these $5$-simplices to form a branched cover of $\overline{\mathcal{M}_{8}}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})$.
@@
  • Initially, $\overline{X}$ is just a copy of the $5$-simplex $\overline{\Delta_{1}}$.
  • Suppose that $\overline{X}$ has a boundary face that is the face $\sigma_{i}(1)$ of some $5$-simplex $\overline{\Delta}$. By Lemma 4.2 and symmetry, we can construct a copy of some $5$-simplex $\overline{\Delta^{\prime}}$ so that $\overline{\Delta}$ and $\overline{\Delta^{\prime}}$ are on different sides of $\sigma_{i}(1)$. Then we glue this copy of $\overline{\Delta^{\prime}}$ to $\overline{X}$ along $\sigma_{i}(1)$.
  • We repeat the process above so that every face in $\overline{X}$ is incident to exactly two $5$-simplices, one on each side.
Let $X_{2}$ be the union of the codimension-two boundaries of the $5$-simplices in $\overline{X}$. Since every face in $\overline{X}$ is incident to exactly two $5$-simplices, $\overline{X}\setminus X_{2}$ is a manifold. There is a continuous map $\iota\mathrel{\mathop{\ordinarycolon}}\overline{X}\rightarrow\overline{\mathcal{M}_{8}}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})$ such that, restricted to every $5$-simplex in $\overline{X}$, $\iota$ is the inclusion map by viewing the $5$-simplex as a subset of $\overline{\mathcal{M}_{8}}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})$. Although $\overline{X}$ contains infinitely many $5$-simplices, there are finitely many loop arrangements involved, so $\iota(\overline{X})$ is a finite union of closed sets. Hence, $\iota(\overline{X})$ is closed in $\overline{\mathcal{M}_{8}}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})$, which implies that $\iota(\overline{X})\setminus\iota(X_{2})$ is closed in $\overline{\mathcal{M}_{8}}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})\setminus\iota(X_{2})$.
@@
Thus, every surface in $\overline{\mathcal{M}_{8}}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})\setminus\iota(X_{2})$ has a parallelogram decomposition. By definition of $X_{2}$, every surface in $\iota(X_{2})$ has a parallelogram decomposition as well. The decomposition is invariant under the antipodal map by our construction. In addition, there are $6$ loops in each loop arrangement, every two of which meet at two antipodal points since they are great circles. In conclusion, we have proved the following result:
@@
For instance, consider the space of centrally symmetric polyhedral surfaces with $8$ unlabeled vertices, whose cone-deficits all equal to $\frac{\pi}{2}$. In this unlabeled space, two surfaces are equivalent if they differ by a Euclidean isometry. Denote its unlabeled moduli space by $\mathcal{M}_{8}(\frac{\pi}{2})$, and its metric completion by $\overline{\mathcal{M}_{8}}(\frac{\pi}{2})$. The following result is a direct consequence of Lemma 4.3:
@@ 5 Polyhedra with $10$ Vertices
In this section, we study the space of centrally symmetric convex polyhedral surfaces with $10$ vertices with prescribed cone-deficits $\delta_{1}$, $\delta_{2}$, $\delta_{3}$, $\delta_{4}$ and $\delta_{5}$. Analogous to the last section, the goal of this section is to show that every surface in $\mathcal{C}_{10}(\delta_{1},\delta_{2},\delta_{3},\delta_{4},\delta_{5})$ can be decomposed into parallelograms. Restricted to this section, a loop arrangement always has $8$ loops labeled by $l_{a}$, $l_{b}$, $l_{c}$, $\dots$, $l_{g}$, $l_{h}$ on a sphere with $8$ vertices labeled by $\pm 1$, $\pm 2$, $\dots$, $\pm 5$.
@@
Thus, $\psi_{1}$ is continuous and invertible. Since $\psi_{1}$ is the composition of $F_{1}$ with another map, $F_{1}$ is continuous and injective on $(0,\infty)^{8}$. As $\mathcal{O}_{1}=F_{1}((0,\infty)^{8})$ and ${\mathcal{C}_{10}}(\delta_{1},\delta_{2}.\delta_{3},\delta_{4},\delta_{5})$ has dimension $8$, the result follows from the invariance of domain. ∎
@@
Without loss of generality, we can assume that $\mathcal{O}$ is $\mathcal{O}_{2}$. Observe from Figure 22 that there are some symmetries among loops in $\mathscr{A}_{2}$. A reflection about the axis on the left interchanges $l_{a}$ and $l_{b}$, $l_{c}$ and $l_{g}$, $l_{d}$ and $l_{h}$. A reflection about the axis on the right interchanges $l_{e}$ and $l_{f}$, $l_{c}$ and $l_{h}$, $l_{d}$ and $l_{g}$. Thus, in summary, we can divide the $8$ loops into three symmetry classes: 1) $l_{a}$, $l_{b}$; 2) $l_{e}$ and $l_{f}$; 3) $l_{c}$, $l_{d}$, $l_{g}$ and $l_{h}$. Therefore, it suffices to discuss the cases when the face $\sigma$ is $\sigma_{b}$, $\sigma_{e}$ or $\sigma_{h}$. Since we have shown that $\overline{\mathcal{O}_{1}}$ and $\overline{\mathcal{O}_{2}}$ are on different sides of $\sigma_{b}$, we only need to discuss the cases $\sigma_{e}$ and $\sigma_{h}$.
@@
There is a Type $4$ space whose closure is on the other side of $\sigma_{d}$ with respect to $\overline{\mathcal{O}_{3}}$.
@@
We have thus finished the discussion with Type $3$ spaces.
@@
Analogous to Section 4, we consider the moduli space $\Delta$ of $\mathcal{O}$, the space of surfaces in $\mathcal{O}$ with area $1$. Let $\overline{\Delta}$ be the completion (with respect to the real hyperbolic metric) of $\Delta$. When $\mathcal{O}$ has Type 1, 2, 3 or 4, a similar argument to Lemma 4.3 shows that $\overline{\Delta}$ is a real hyperbolic ideal $7$-simplex.
@@
The idea of proof is essentially the same as that of Theorem 4.5. We construct infinitely many copies of the moduli spaces of Type 1, 2, 3 or 4. Then we glue their metric completions (which are $7$-simplices) to form a space $\overline{X}$ such that every face belongs to exactly two $7$-simplices, one on each side. By Lemma 5.14, such $\overline{X}$ exists. Then we define the map $\iota$ analogously. Away from the codimension-two boundaries of $\overline{X}$, the openness of $\iota$ is guaranteed by Corollary 5.15. Finally, there are $8$ loops in every loop arrangement, each two of which intersect exactly twice as great circles, hence the number $2\binom{8}{2}$. ∎
@@ 6 Discussions and Future Work
Let $\delta_{1}$, $\delta_{2}$, $\dots$, $\delta_{N}$ be $N$ positive numbers that sum up to $2\pi$. Then every surface in $\overline{\mathcal{M}_{2N}}(\delta_{1},\delta_{2},\dots,\delta_{N})$ can be decomposed into at most $2\binom{2N-2}{2}$ parallelograms, and the decomposition is invariant under the antipodal map.

### 30 Mar 2026 v2

@@ 4 Polyhedra with 88 Vertices
Therefore, $\psi_{1}$ is continuous and invertible. Since $\psi_{1}$ is a composition of $F_{1}$ by another map, $F_{1}$ is injective on $(0,\infty)^{6}$. By Theorem 2.1, $\mathcal{C}_{8}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})$ has real dimension 66. By the invariance of domain, $\mathcal{O}_{1}=F_{1}((0,\infty)^{6})$ is open in $\mathcal{C}_{8}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})$. ∎
@@
Given a space $\mathcal{O}$ arising from a loop arrangement $\mathscr{A}$, we can consider the space of polyhedral surfaces in $\mathcal{O}$ with surface area 11. We call this space the moduli space arising from $\mathscr{A}$, and denote it by $\Delta$. Let $\overline{\Delta}\subseteq\overline{\mathcal{M}_{8}}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})$ be the metric completion of $\Delta$ with respect to the real hyperbolic metric in Theorem $2.$ We denote a face of $\overline{\Delta}$ by $\sigma_{i}({1})$ (where 11 refers to the surface area) if it is a subset of the face $\sigma_{i}$ in $\overline{\mathcal{O}}$. If $\mathscr{A}^{\prime}$ is another loop arrangement that differs from $\mathscr{A}$ by $l_{i}$ only, we say that $\overline{\Delta}$ and $\overline{\Delta^{\prime}}$ are on different sides of $\bm{\sigma_{i}({1})}$ if and only if $\overline{\mathcal{O}}$ and $\overline{\mathcal{O}^{\prime}}$ are on different sides of $\sigma_{i}$.
@@
The space $\overline{\Delta_{1}}$ has the structure of a real hyperbolic regular ideal 55-simplex.
@@
Combinatorically, $\overline{\Delta_{1}}$ is a polytope with six vertices and six facets. Each vertex of $\overline{\Delta_{1}}$ has five vanishing coordinates among six with respect to $\mathscr{A}_{1}$. This corresponds to the case where the surface degenerates to a line. Note that each vertex of $\overline{\Delta_{1}}$ belongs to five facets of $\overline{\Delta_{1}}$, and each facet contains five vertices. Therefore, $\overline{\Delta_{1}}$ is a 55-simplex by Proposition 2.16 in [13].
@@
Similar to $\mathscr{A}_{2}$, by permuting the the labels of vertices and loops in $\mathscr{A}_{1}$, we can obtain more loop arrangements that are different from but combinatorically equivalent to $\mathscr{A}_{1}$. It is clear that there are finitely many such loop arrangements. We consider the metric completions of the moduli spaces arising from these loop arrangements. For convenience, we will call them “55-simplices” due to Lemma 4.3. Later, we are going to construct infinitely many copies of such 55-simplices and regard them as distinct spaces, though some of them arise from the same loop arrangement. Our goal is to glue these 55-simplices to form a branched cover of $\overline{\mathcal{M}_{8}}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})$.
@@
  • Initially, $\overline{X}$ is just a copy of the 55-simplex $\overline{\Delta_{1}}$.
  • Suppose that $\overline{X}$ has a boundary face that is the face $\sigma_{i}(1)$ of some 55-simplex $\overline{\Delta}$. By Lemma 4.2 and symmetry, we can construct a copy of some 55-simplex $\overline{\Delta^{\prime}}$ so that $\overline{\Delta}$ and $\overline{\Delta^{\prime}}$ are on different sides of $\sigma_{i}(1)$. Then we glue this copy of $\overline{\Delta^{\prime}}$ to $\overline{X}$ along $\sigma_{i}(1)$.
  • We repeat the process above so that every face in $\overline{X}$ is incident to exactly two 55-simplices, one on each side.
Let $X_{2}$ be the union of the codimension-two boundaries of the 55-simplices in $\overline{X}$. Since every face in $\overline{X}$ is incident to exactly two 55-simplices, $\overline{X}\setminus X_{2}$ is a manifold. There is a continuous map $\iota\mathrel{\mathop{\ordinarycolon}}\overline{X}\rightarrow\overline{\mathcal{M}_{8}}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})$ such that, restricted to every 55-simplex in $\overline{X}$, $\iota$ is the inclusion map by viewing the 55-simplex as a subset of $\overline{\mathcal{M}_{8}}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})$. Although $\overline{X}$ contains infinitely many 55-simplices, there are finitely many loop arrangements involved, so $\iota(\overline{X})$ is a finite union of closed sets. Hence, $\iota(\overline{X})$ is closed in $\overline{\mathcal{M}_{8}}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})$, which implies that $\iota(\overline{X})\setminus\iota(X_{2})$ is closed in $\overline{\mathcal{M}_{8}}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})\setminus\iota(X_{2})$.
@@
Thus, every surface in $\overline{\mathcal{M}_{8}}(\delta_{1},\delta_{2},\delta_{3},\delta_{4})\setminus\iota(X_{2})$ has a parallelogram decomposition. By definition of $X_{2}$, every surface in $\iota(X_{2})$ has a parallelogram decomposition as well. The decomposition is invariant under the antipodal map by our construction. In addition, there are 66 loops in each loop arrangement, every two of which meet at two antipodal points since they are great circles. In conclusion, we have proved the following result:
@@
For instance, consider the space of centrally symmetric polyhedral surfaces with 88 unlabeled vertices, whose cone-deficits all equal to $\frac{\pi}{2}$. In this unlabeled space, two surfaces are equivalent if they differ by a Euclidean isometry. Denote its unlabeled moduli space by $\mathcal{M}_{8}(\frac{\pi}{2})$, and its metric completion by $\overline{\mathcal{M}_{8}}(\frac{\pi}{2})$. The following result is a direct consequence of Lemma 4.3:
@@ 5 Polyhedra with $10$ Vertices
In this section, we study the space of centrally symmetric convex polyhedral surfaces with $10$ vertices with prescribed cone-deficits $\delta_{1}$, $\delta_{2}$, $\delta_{3}$, $\delta_{4}$ and $\delta_{5}$. Analogous to the last section, the goal of this section is to show that every surface in $\mathcal{C}_{10}(\delta_{1},\delta_{2},\delta_{3},\delta_{4},\delta_{5})$ can be decomposed into parallelograms. Restricted to this section, a loop arrangement always has 88 loops labeled by $l_{a}$, $l_{b}$, $l_{c}$, $\dots$, $l_{g}$, $l_{h}$ on a sphere with 88 vertices labeled by $\pm 1$, $\pm 2$, $\dots$, $\pm 5$.
@@
Thus, $\psi_{1}$ is continuous and invertible. Since $\psi_{1}$ is the composition of $F_{1}$ with another map, $F_{1}$ is continuous and injective on $(0,\infty)^{8}$. As $\mathcal{O}_{1}=F_{1}((0,\infty)^{8})$ and ${\mathcal{C}_{10}}(\delta_{1},\delta_{2}.\delta_{3},\delta_{4},\delta_{5})$ has dimension 88, the result follows from the invariance of domain. ∎
@@
Without loss of generality, we can assume that $\mathcal{O}$ is $\mathcal{O}_{2}$. Observe from Figure 22 that there are some symmetries among loops in $\mathscr{A}_{2}$. A reflection about the axis on the left interchanges $l_{a}$ and $l_{b}$, $l_{c}$ and $l_{g}$, $l_{d}$ and $l_{h}$. A reflection about the axis on the right interchanges $l_{e}$ and $l_{f}$, $l_{c}$ and $l_{h}$, $l_{d}$ and $l_{g}$. Thus, in summary, we can divide the 88 loops into three symmetry classes: 1) $l_{a}$, $l_{b}$; 2) $l_{e}$ and $l_{f}$; 3) $l_{c}$, $l_{d}$, $l_{g}$ and $l_{h}$. Therefore, it suffices to discuss the cases when the face $\sigma$ is $\sigma_{b}$, $\sigma_{e}$ or $\sigma_{h}$. Since we have shown that $\overline{\mathcal{O}_{1}}$ and $\overline{\mathcal{O}_{2}}$ are on different sides of $\sigma_{b}$, we only need to discuss the cases $\sigma_{e}$ and $\sigma_{h}$.
@@
There is a Type 44 space whose closure is on the other side of $\sigma_{d}$ with respect to $\overline{\mathcal{O}_{3}}$.
@@
We have thus finished the discussion with Type 33 spaces.
@@
Analogous to Section 4, we consider the moduli space $\Delta$ of $\mathcal{O}$, the space of surfaces in $\mathcal{O}$ with area 11. Let $\overline{\Delta}$ be the completion (with respect to the real hyperbolic metric) of $\Delta$. When $\mathcal{O}$ has Type 1, 2, 3 or 4, a similar argument to Lemma 4.3 shows that $\overline{\Delta}$ is a real hyperbolic ideal 77-simplex.
@@
The idea of proof is essentially the same as that of Theorem 4.5. We construct infinitely many copies of the moduli spaces of Type 1, 2, 3 or 4. Then we glue their metric completions (which are 77-simplices) to form a space $\overline{X}$ such that every face belongs to exactly two 77-simplices, one on each side. By Lemma 5.14, such $\overline{X}$ exists. Then we define the map $\iota$ analogously. Away from the codimension-two boundaries of $\overline{X}$, the openness of $\iota$ is guaranteed by Corollary 5.15. Finally, there are 88 loops in every loop arrangement, each two of which intersect exactly twice as great circles, hence the number $2\binom{8}{2}$. ∎
@@ 6 Discussions and Future Work
Let $\delta_{1}$, $\delta_{2}$, $\dots$, $\delta_{N}$ be NN positive numbers that sum up to $2\pi$. Then every surface in $\overline{\mathcal{M}_{2N}}(\delta_{1},\delta_{2},\dots,\delta_{N})$ can be decomposed into at most $2\binom{2N-2}{2}$ parallelograms, and the decomposition is invariant under the antipodal map.